# Towards the Causal Complete Cause of Multi-Modal Representation Learning

### v4

@@ 3 Causal Complete Cause
To access the learning of causal sufficient and necessary representations, in this section, we first provide the definition of causal sufficiency and necessity in MML, i.e., Causal Complete Cause ($C^{3}$). Next, we discuss the identifiability of $C^{3}$, which ensures the quantification of $C^{3}$ through observable data in practice, even in corner cases (non-monotonicity and non-exogeneity). Finally, we give the measurement of $C^{3}$, i.e., $C^{3}$ risk to measure the probability of whether the learned MML representation is causally complete.
@@ 3.1 Definition of $C^{3}$
Definition 3.1 indicates that when Z with a high $C^{3}$ score, it has a high probability of being the causal complete cause of Y. The two terms in Eq.1 correspond to sufficiency and necessity, respectively. Specifically, the sufficiency term $P(\text{Y}_{do(\text{Z}=c)}=y\mid\text{Z}=\bar{c},\text{Y}\not=y)$ means that even when $\text{Z}=\bar{c}$ and $\text{Y}\neq y$, intervening to set $\text{Z}=c$ results in a high probability of $\text{Y}=y$, indicating that $\text{Z}=c$ has a sufficient causal effect on the occurrence of $\text{Y}=y$. Correspondingly, for necessity term $P(\text{Y}_{do(\text{Z}=\bar{c})}\not=y\mid\text{Z}=c,\text{Y}=y)$, even when $\text{Z}=c$ and $\text{Y}=y$, intervening to set $\text{Z}=\bar{c}$ results in $\text{Y}\neq y$, indicating that $\text{Z}=c$ is necessary for the occurrence of $\text{Y}=y$. Under this condition, the learned representation can be divided into sufficient but unnecessary causes, necessary but insufficient causes, sufficient and necessary causes, and insufficient and unnecessary causes (Appendix D for detailed analyses). We aim to constrain the $C^{3}$ score of the learned representation, i.e., extract sufficient and necessary causes, to achieve robust and accurate MML.
3.2 Identifiability of $C^{3}$
@@
Since it is difficult to obtain all samples in the multi-modal data distribution, especially in real systems, e.g., the counterfactual data in the definition of $C^{3}$ is difficult to obtain (Kusner et al., 2017; Morgan & Winship, 2015), calculating the probability of $C^{3}$ is still is a challenging issue. To access the calculation of $C^{3}$ based on observable MML data, we discuss the identifiability of $C^{3}$ in this section.
@@
Identifiability refers to the ability to uniquely infer causal effects from observable data under given assumptions (Pearl, 2009). For MML settings, it means the causal factors within the representation $Z$ can uniquely determine the label $Y$ from the sample $X$. To ensure reliable estimation, we constrain the $C^{3}$ score of the learned representation $Z$ extracted from $X$, ensuring its identifiability. In previous studies (Tian & Pearl, 2002; Yang et al., 2024), identifying causal probabilities typically assumes that statistical data are derived under exogeneity and monotonicity conditions (Definition D.1 and Definition D.2 in Appendix D.4). Exogeneity refers to where the external intervention’s impact on the conditional distribution $P(\text{Y}|\text{Z})$ is negligible when the variable Z is exogenous to Y. This ensures the learned Z to recognize $\text{F}_{c}$ without being affected by $\text{F}_{s}$. Monotonicity, on the other hand, describes the consistent, unidirectional effect of the representation Z on Y. This ensures that the impact of the learned representation Z on the labeling decision will not be reversed as other factors change. However, in practice, non-trivial cases often arise: the inseparability of generating factors in the latent space leads to spurious correlations, i.e., the learned representations contain non-causal $\text{F}_{s}$ (Figure 2); Meanwhile, modality conflicts and the high-dimensional nonlinear interactions (Wang et al., 2023d; Huang et al., 2021b) undermine monotonicity. These challenges render traditional identifiability conditions inapplicable. To address this, we relax the above assumptions and explore the identifiability of $C^{3}$, accessing the calculation of $C^{3}$ score.
@@
Consider the effects of spurious correlations, i.e., $P(\text{Y}_{do(\text{Z}=c)})\neq P(\text{Y}\mid\text{Z}=c)$, we first relax the exogeneity assumption. Based on Definition 3.1, we extend Theorem 9.2.15 in (Pearl, 2009) to MML for $C^{3}$, obtaining:
Theorem 3.2 (Identifiability under Non-Exogeneity).
@@
Theorem 3.2 establishes that, under the monotonicity condition, $C^{3}$ can be estimated using observable multi-modal data, thereby quantifying $C^{3}$. Specifically, by modeling the effects of spurious correlations ($\text{F}_{s}$ on Y) as non-causal pathways in the SCM, interventions with the $do$-operator, such as $P(\text{Y}_{do(\text{Z}=c)})$ and $P(\text{Y}_{do(\text{Z}=\bar{c})})$, facilitates reliable causal effect estimation. For counterfactual terms ($\text{Z}=\bar{c}$) involved in the necessity probability, the model $f_{\theta}$ imposes locally reversible constraints to ensure identifiability, as discussed in (Galles & Pearl, 1998). Under these conditions, the probability of causal sufficiency $C^{3}_{su}$ and the probability of causal necessity $C^{3}_{ne}$ can be estimated as follows:
@@
Theorem 3.3 (Identifiability under Non-Monotonicity and Non-Exogeneity).
@@ 3.3 Measurement of $C^{3}$
In Theorem 3.4, we propose using the self-attention mechanism to model the instrumental variable $V$ for $C^{3}$. Specifically, the alignment scores $s_{ij}$ is to capture inter-modal interactions for label. Then, by strengthening the correlation between $V$ and Z, the learned representation Z is guided to approach $\text{F}_{c}$ while weakening the correlation with $\text{F}_{s}$. Furthermore, consider inter-modal misalignment, $\|q_{i}-k_{j}\|^{2}$ with weight $\alpha$ is introduced as a distance-based adjustment to diminish the impact of modalities with large feature distances on attention weights. It satisfies three conditions: (i) Relevance: $V$ manipulate $\text{F}_{c}$; (ii) Independence: $V$ is conditionally independent from $\text{F}_{s}$; and (iii) Exclusion: the influence of $V$ on Y is completely indirectly realized through $\text{F}_{c}$. The detailed proofs are provided in Appendix A.3.
@@ 4 Performance Guarantee with $C^{3}$ Risk
Theorem 4.1 derives the upper bound on the generalization error of $C^{3}$ risk, relating its empirical error on the training data to its performance on unseen samples. Specifically, the generalization error is bounded by the sum of the empirical error, the Rademacher complexity of the hypothesis class of linear classifier, and a confidence term that controls the probability of exceeding the bound. This result provides a theoretical performance guarantee for optimizing MML models with $C^{3}$ risk. See Appendix A.5 for detailed proofs.
5 Learning Causal Complete Representations
6 Experiments
Table 1: Performance comparison when 50% samples are corrupted with Gaussian noise i.e., zero mean with the variance of $N$. “(N, Avg.)” and “(N, Worst.)” denotes the average and worst-case accuracy. The best results are highlighted in bold.
@@ 6.1 Experimental Settings
Implementation Details We use a three-layer MLP with activation functions (Clevert et al., 2015) as the representation learner. The hidden vector dimensions of each layer are specified as 64, 32, and 128, while the learned representation is 64. For optimization, we employ the Adam optimizer (Kingma & Ba, 2014) with Momentum and weight decay set at $0.8$ and $10^{-4}$. The initial learning rate is established at 0.1, with the flexibility for linear scaling as required. Additionally, we use grid search to set the hyperparameters $\lambda_{v}=0.75$ and $\lambda_{fe}=0.4$. All experimental procedures are executed in five runs via NVIDIA RTX A6000 GPUs.
@@ 6.2 Results
Performance and robustness analysis
To evaluate the effectiveness of $C^{3}$, we record the average and worst-case accuracy of various MML baselines and introduce $C^{3}$ under Gaussian noise (for image modality) and blank noise (for text modality) following (Han et al., 2020; Zhang et al., 2023d; Ma et al., 2021). The results are shown in Table 1. We can observe that $C^{3}$ achieves stable improvements in both the average and worst-case accuracy. This proves the superior effectiveness and robustness of $C^{3}$.
When faces the problem of missing modalities
Considering the missing modality issues, we evaluate the performance of $C^{3}$ and several strong baselines (Wang et al., 2023b; Bakas et al., 2018; Zhang et al., 2022) on all 15 possible combinations of missing modalities on BraTS. From Table 2, we can observe that (i) $C^{3}$ brings significant performance improvements; (ii) $C^{3}$ can reduce the learning gap for the representations on different modal semantics, i.e., reducing the accuracy gap of learning on the difficult-to-identify Fl and T1 modalities and the easy T1c. This demonstrates the superiority of $C^{3}$ and the advantage of causally complete representation in missing modality issues.
Learning causal complete representations
To evaluate $C^{3}$’s ability to extract causal complete causes, we conduct experiments: (i) construct four types of MML data, i.e., sufficient and necessary (SNC), sufficient but unnecessary (SC), necessary but insufficient causes (NC), and spurious correlations (SP) following (Yang et al., 2024) (see Appendix H for more details); then (ii) evaluate their correlation with the learned representation based on the distance metric (Jones et al., 1995), i.e., the higher the score, the stronger the correlation. The results are shown in Figure 4 (see Appendix D and H for details). The representations learned by $C^{3}$ have a higher correlation with SNC and lower with SP. This proves that $C^{3}$ can learn causal complete representations more effectively while other methods are hard to achieve.
Ablation Study
To evaluate the effect of each item in $C^{3}$ (Eq.10), we evaluate the performance change on LCKD+$C^{3}$. We consider the standard and corner case (missing modality), and conduct experiments on FOOD 101 and BraTS, as shown in Figure 4. We can observe that each item plays an important role. See Appendix H for more experiments, e.g., model efficiency and parameter sensitivity.
@@ 7 Related Work
Multi-modal learning aims to learn good representations through multiple modalities for accurate prediction. Recently, multiple methods (Xu et al., 2023; Jia et al., 2022b; Wang et al., 2023c; Fan et al., 2024) have been proposed to solve MML tasks through modality consistency, e.g., tokenize diverse modalities into sequences and utilize Transformers for joint learning (Bao et al., 2022; Wang et al., 2022), whereas CLIP (Radford et al., 2021; Fan et al., 2024), ALIGN (Jia et al., 2021), etc. employ distinct encoders for each modality and utilize contrastive loss to synchronize features. These methods align features from different modalities into the same space to capture the primary events. Another type of work (Jiang et al., 2023; Dong et al., 2024; Wang et al., 2023a; Wu et al., 2020) proposed to additionally focus on modality specificity. They divide the MML features into modality-shared and modality-specific components and learn separately. However, all these methods may result in learning insufficient or unnecessary information. Meanwhile, they rely on strong assumptions like semantic alignment (Akbari et al., 2021; Wang et al., 2023c; Lu et al., 2022; Zhang et al., 2023b) and may limited to ideal data (Ge et al., 2023; Zhang et al., 2023a), affecting model performance. For causal sufficiency and necessity, (Pearl, 2009) first proposed the related concepts, and (Yang et al., 2024) applied it to domain generalization with assumptions in an ideal environment. The differences between our methodology and theirs include problem settings, theorems, implementations, learning objectives, etc. (See Appendix D.6 for detailed comparison). We explore the MML-specific concepts of causal completeness without strong assumptions, ensuring the effectiveness of the learned representations.

### v6

@@ 3 Causal Complete Cause
To access the learning of causal sufficient and necessary representations, in this section, we first provide the definition of causal sufficiency and necessity in MML, i.e., Causal Complete Cause ($C^{3}$). Next, we discuss the identifiability of $C^{3}$, which ensures the quantification of $C^{3}$ through observable data in practice, even in corner cases (non-monotonicity and non-exogeneity). Finally, we give the measurement of $C^{3}$, i.e., $C^{3}$ risk, to measure the probability of whether the learned MML representation is causally complete.
@@ 3.1 Definition of $C^{3}$
Definition 3.1 indicates that when Z with a high $C^{3}$ score, it has a high probability of being the causal complete cause of Y. The two terms in Eq.1 correspond to sufficiency and necessity, respectively. Specifically, the sufficiency term $P(\text{Y}_{do(\text{Z}=c)}=y\mid\text{Z}=\bar{c},\text{Y}\not=y)$ means that even when $\text{Z}=\bar{c}$ and $\text{Y}\neq y$, intervening to set $\text{Z}=c$ results in a high probability of $\text{Y}=y$, indicating that $\text{Z}=c$ has a sufficient causal effect on the occurrence of $\text{Y}=y$. Correspondingly, for necessity term $P(\text{Y}_{do(\text{Z}=\bar{c})}\not=y\mid\text{Z}=c,\text{Y}=y)$, even when $\text{Z}=c$ and $\text{Y}=y$, intervening to set $\text{Z}=\bar{c}$ results in $\text{Y}\neq y$, indicating that $\text{Z}=c$ is necessary for the occurrence of $\text{Y}=y$. Under this condition, the learned representation can be divided into sufficient but unnecessary causes, necessary but insufficient causes, sufficient and necessary causes, and insufficient and unnecessary causes (Appendix D.1 for detailed analyses). We aim to constrain the $C^{3}$ score of the learned representation, i.e., extract sufficient and necessary causes, to achieve robust and accurate MML.
3.2 Causal Identifiability of $C^{3}$
@@
Since it is difficult to obtain all samples in the multi-modal data distribution, especially in real systems, e.g., the counterfactual data in the definition of $C^{3}$ is difficult to obtain (Kusner et al., 2017; Morgan & Winship, 2015), calculating the probability of $C^{3}$ is still a challenging issue. To access the calculation of $C^{3}$ based on observable MML data, we discuss the causal identifiability of $C^{3}$ in this section.
@@
Identifiability refers to the ability to uniquely infer causal effects from observable data under given assumptions (Pearl, 2009). For MML settings, it means the causal factors within the representation $Z$ can uniquely determine the label $Y$ from the sample $X$. To ensure reliable estimation, we constrain the $C^{3}$ score of the learned representation $Z$ extracted from $X$, ensuring its identifiability. In previous studies (Tian & Pearl, 2002; Yang et al., 2024), identifying causal probabilities typically assumes that statistical data are derived under exogeneity and monotonicity conditions (Definition D.1 and Definition D.2 in Appendix D.3). Exogeneity refers to where the external intervention’s impact on the conditional distribution $P(\text{Y}|\text{Z})$ is negligible when the variable Z is exogenous to Y. This ensures the learned Z to recognize $\text{F}_{c}$ without being affected by $\text{F}_{s}$. Monotonicity, on the other hand, describes the consistent, unidirectional effect of the representation Z on Y. This ensures that the impact of the learned Z on the labeling decision will not be reversed as other factors change. However, in practice, non-trivial cases often arise: the inseparability of generating factors in the latent space leads to spurious correlations, i.e., the learned representations Z contain non-causal $\text{F}_{s}$ (Figure 2); Meanwhile, modality conflicts and the high-dimensional nonlinear interactions (Wang et al., 2024; Huang et al., 2021b) undermine monotonicity. These challenges render traditional identifiability conditions inapplicable. To address this, we relax the above assumptions and explore the identifiability of $C^{3}$, assessing the calculation of the $C^{3}$ score.
@@
Consider the effects of spurious correlations, i.e., $P(\text{Y}_{do(\text{Z}=c)})\neq P(\text{Y}\mid\text{Z}=c)$, we first relax the exogeneity assumption (with definition and analysis in Appendix D.3). Based on Definition 3.1, we extend Theorem 9.2.15 in (Pearl, 2009) to MML for $C^{3}$, obtaining:
Theorem 3.2 (Causal Identifiability under Non-Exogeneity).
@@
Theorem 3.2 establishes that, under the monotonicity condition, $C^{3}$ can be estimated using observable multi-modal data, thereby quantifying $C^{3}$. The local invertibility (Proposition D.3) states that the model can uniquely recover the distribution of $s$ from the conditional distribution of $Y$ given its parents $\mathrm{Pa}(Y)$. Specifically, by modeling the effects of spurious correlations ($\text{F}_{s}$ on Y) as non-causal pathways in the SCM, interventions with the $do$-operator, such as $P(\text{Y}_{do(\text{Z}=c)})$ and $P(\text{Y}_{do(\text{Z}=\bar{c})})$, facilitates reliable causal effect estimation. For counterfactual terms ($\text{Z}=\bar{c}$) involved in the necessity probability, the model $f_{\theta}$ imposes locally reversible constraints to ensure identifiability, as discussed in (Galles & Pearl, 1998). Under these conditions, the probability of causal sufficiency $C^{3}_{su}$ and the probability of causal necessity $C^{3}_{ne}$ can be estimated as follows:
@@
Theorem 3.3 (Causal Identifiability under Non-Monotonicity and Non-Exogeneity).
@@ 3.3 Measurement of $C^{3}$
In Theorem 3.4, we propose using the self-attention mechanism to model the instrumental variable $V$ for $C^{3}$. Specifically, the alignment score $s_{ij}$ is to capture inter-modal interactions for labels. Then, by strengthening the correlation between $V$ and Z, the learned representation Z is guided to approach $\text{F}_{c}$ while weakening the correlation with $\text{F}_{s}$. Furthermore, consider inter-modal misalignment, $\|q_{i}-k_{j}\|^{2}$ with weight $\alpha$ is introduced as a distance-based adjustment to diminish the impact of modalities with large feature distances on attention weights. It satisfies three conditions: (i) Relevance: $V$ manipulate $\text{F}_{c}$; (ii) Independence: $V$ is conditionally independent from $\text{F}_{s}$; and (iii) Exclusion: the influence of $V$ on Y is completely indirectly realized through $\text{F}_{c}$. The detailed proofs are provided in Appendix A.3.
@@ 4 Performance Guarantee with $C^{3}$ Risk
Theorem 4.1 derives the upper bound on the generalization error of $C^{3}$ risk, relating its empirical error on the training data to its performance on unseen samples. Specifically, the generalization error is bounded by the sum of the empirical error, the Rademacher complexity of the hypothesis class of the linear classifier, and a confidence term that controls the probability of exceeding the bound. This result provides a theoretical performance guarantee for optimizing MML models with $C^{3}$ risk. See Appendix A.5 for detailed proofs.
Table 1: Performance comparison when 50% samples are corrupted with Gaussian noise, i.e., zero mean with the variance of $N$. “(N, Avg.)” and “(N, Worst.)” denotes the average and worst-case accuracy. The best results are highlighted in bold.
@@ 6.1 Experimental Settings
Implementation Details We use a three-layer MLP with activation functions (Clevert et al., 2016) as the representation learner. The hidden vector dimensions of each layer are specified as 64, 32, and 128, while the learned representation is 64. For optimization, we employ the Adam optimizer (Kingma & Ba, 2015) with Momentum and weight decay set at $0.8$ and $10^{-4}$. The initial learning rate is established at 0.1, with the flexibility for linear scaling as required. Additionally, we use grid search to set the hyperparameters $\lambda_{v}=0.75$ and $\lambda_{fe}=0.4$. All experimental procedures are executed in five runs via NVIDIA RTX A6000 GPUs. Unless otherwise specified, all baselines and corresponding settings are reproduced following the original configurations mentioned in the papers. Codes can be found in Github111 Codes can be found in https://github.com/WangJingyao07/Multi-Modal-Base.
@@ 6.2 Results
Performance and robustness analysis To evaluate the effectiveness of $C^{3}$, we record the performance change across various MML baselines after introducing $C^{3}$ under Gaussian noise (for image modality) and blank noise (for text modality) following (Han et al., 2020; Zhang et al., 2023d; Ma et al., 2021). We record both the average and worst-case accuracy. The results are shown in Table 1. We can observe that $C^{3}$ achieves stable improvements in both the average and worst-case accuracy. This proves the superior effectiveness and robustness of $C^{3}$.
When faces the problem of missing modalities Considering the missing modality issues, we evaluate the performance of $C^{3}$ and several strong baselines (Wang et al., 2023b; Bakas et al., 2018; Zhang et al., 2022b) on all 15 possible combinations of missing modalities on BraTS. From Table 2, we can observe that (i) $C^{3}$ brings significant performance improvements; (ii) $C^{3}$ can reduce the learning gap for the representations on different modal semantics, i.e., reducing the accuracy gap of learning on the difficult-to-identify Fl and T1 modalities and the easy T1c. This demonstrates the superiority of $C^{3}$ and the advantage of causally complete representation in missing modality issues.
Learning causal complete representations To evaluate $C^{3}$’s ability to extract causal complete causes, we conduct experiments: (i) construct four types of MML data, i.e., sufficient and necessary (SNC), sufficient but unnecessary (SC), necessary but insufficient causes (NC), and spurious correlations (SP) following (Yang et al., 2024) (see Appendix H for more details); then (ii) evaluate their correlation with the learned representation based on the distance metric (Jones et al., 1995), i.e., the higher the score, the stronger the correlation. The results are shown in Figure 4 (see Appendix D and H for details). The representations learned by $C^{3}$ have a higher correlation with SNC and lower with SP. This proves that $C^{3}$ can learn causal complete representations more effectively, while other methods are hard to achieve.
Ablation Study To evaluate the effect of each item in $C^{3}$ (Eq.10), we evaluate the performance change on LCKD+$C^{3}$. We record the performance changes before and after the model introduced specific terms. We consider the standard and corner case (missing modality), and conduct experiments on FOOD 101 and BraTS. The results are shown in Figure 4. We can observe that each item plays an important role. See Appendix H for more experiments, including the analyses about model efficiency and parameter sensitivity. We provide a brief summary about all the experiments conducted in this paper, as shown in Table 3.
@@ 7 Related Work
Multi-modal learning aims to learn good representations through multiple modalities for accurate prediction. Recently, multiple methods (Xu et al., 2023; Jia et al., 2022b; Wang et al., 2023c; Fan et al., 2024) have been proposed to solve MML tasks through modality consistency, e.g., tokenize diverse modalities into sequences and utilize Transformers for joint learning (Bao et al., 2022; Wang et al., 2022), whereas CLIP (Radford et al., 2021; Fan et al., 2024), ALIGN (Jia et al., 2021), etc. employ distinct encoders for each modality and utilize contrastive loss to synchronize features. These methods align features from different modalities into the same space to capture the primary events. Another type of work (Jiang et al., 2023; Dong et al., 2024; Wang et al., 2023a; Wu et al., 2020) proposed to additionally focus on modality specificity. They divide the MML features into modality-shared and modality-specific components and learn separately. However, all these methods may result in learning insufficient or unnecessary information. Meanwhile, they rely on strong assumptions like semantic alignment (Akbari et al., 2021; Wang et al., 2023c; Lu et al., 2022; Zhang et al., 2023b) and may be limited to ideal data (Ge et al., 2023; Zhang et al., 2023a), affecting model performance. For causal sufficiency and necessity, (Pearl, 2009) first proposed the related concepts, and (Yang et al., 2024) applied it to domain generalization with assumptions in an ideal environment. The differences between our methodology and previous causal-related works include problem settings, theorems, implementations, learning objectives, etc. (More comparison and analyses are provided in Appendix D.6). We explore the MML-specific concepts of causal completeness without strong assumptions, ensuring the effectiveness of the learned representations.